Sifton Electronics Corporation manufactures and assembles electronic motor drives for video cameras. The company assembles the motor drives for several accounts. The process consists of a lean cell for each customer. The following information relates only to one customer's lean cell for the coming year. Projected labor and overhead, $3,839,400; materials costs, $30 per unit. Planned production included 5,184 hours to produce 16,200 motor drives. Actual production for August was 1,890 units, and motor drives shipped amounted to 1,440 units.

From the foregoing information, determine the manufacturing cost per unit.

a.$267.00

b.$237.00

c.$207.00

d.$770.63
Answer #1
Calculation or manufacturing cost per unit
projected labor & overhead per month (3839400/12) = 319950
labor & overhead cost per hour (3839400/5184 ) = 740.625
unit per hour (16200/5184) = 3.125
required time per hour = 1/3.125 = 0.32
manufacturing cost per unit = 740.625*0.32 = 237
Option "b" $237 is Correct